# A functional limit theorem for a dynamical system with an observable maximised on a Cantor set

### v1

@@ 4.4. Dependence requirements
In order to prove that the conditions Д$\D_{q_{n}}$ and Д$\D^{\prime}_{q_{n}}$ from Section 3.2 hold in our setting we will approximate the sets $A_{n,\ell}^{(q_{n})}$ by a union of cylinders of controlled depth and then use the excellent mixing properties on cylinders in the spirit of [HP14, DHY21, BF24], for example.
@@
$U_{n}(\tau^{\prime},\tau^{\prime\prime})$ is made up of holes in the dynamical construction of $\Lambda$ (i.e. subsets of the form $\Lambda_{k}\setminus\Lambda_{k+1}$ where $k\in\mathbb{N}$), where $j_{n,\tau^{\prime}}$ is as in (4.2). Observe also that, for $n$ sufficiently large,
$\begin{split}U_{n}(\tau^{\prime},\tau^{\prime\prime})^{(q_{n})}&=U_{n}(\tau^{\prime},\tau^{\prime\prime})\cap G^{-1}(U_{n}(\tau^{\prime},\tau^{\prime\prime}))^{c}\cap\dots\cap G^{-q_{n}}(U_{n}(\tau^{\prime},\tau^{\prime\prime}))^{c}\\ &=U_{n}(\tau^{\prime},\tau^{\prime\prime})\cap G^{-1}(U_{n}(\tau^{\prime},\tau^{\prime\prime}))^{c}\end{split}$ (4.8)
so, in particular, we can take $q_{n}=1$ for all $n\in\mathbb{N}$.
@@
where $\beta,\gamma\in[0,1]$ and $\theta$ is as in Proposition 4.6 (as $H^{(j_{n,\tau^{\prime}})}$ is a fraction of $\Lambda_{j_{n,\tau^{\prime}}-1}\setminus\Lambda_{j_{n,\tau^{\prime}}}$ while $K$ is a fraction of $\Lambda_{j_{n,\tau^{\prime}}}\setminus\Lambda_{j_{n,\tau^{\prime}}+1}$). Observe that the larger the set $H^{(j_{n,\tau^{\prime}})}$ is, the smaller the set $K$ is and vice-versa. This means that, in particular, $\beta$ and $\gamma$ cannot be simultaneously equal to 00 or to $1$.
On the other hand, taking into consideration the definition of the function $\psi$ and of the approximating cylinders, as well as the fact that $G$ is a full branched Markov linear map (which, in particular, means that there exist $\tilde{C}>0$ and $10<\eta<$ 1 such that $\mathfrak{m}(\omega)\leq\tilde{C}\eta^{n}$, for all $n\in\mathbb{N}$ and $\omega\in\mathcal{P}^{n}$), we claim that
@@
As a result, we only need to check that the conditions Д$\D_{q_{n}}^{*}$ and Д$\D^{{}^{\prime}*}_{q_{n}}$ are verified for the cylinder approximating sets. To this end, we next define a strong form of mixing, which is a property of the maps we are considering here.
@@
We remark that $G$ is exponentially $\phi$-mixing, see for example [AN05, Theorem 1]. Recall that, as observed above, we only need to check the conditions Д$\D_{q_{n}}^{*}$, Д$\D_{q_{n}}^{{}^{\prime}*}$, where the set $A_{n,\ell}^{(q_{n})}$ is replaced by the approximations $\Gamma_{n}^{+}$ and $\Gamma_{n}^{-}$. In what follows, we will only check such conditions for $\Gamma_{n}^{+}$, but the argument applies with minor adjustments to $\Gamma_{n}^{-}$.
Proof of Condition Д$\D_{q_{n}}^{*}$.
@@
Proof of Condition Д$\D_{q_{n}}^{{}^{\prime}*}$.
@@
for some $\tilde{C},\hat{C}>0$. The second term goes to 00 by (3.2), Lemma 4.8 and the fact that $k_{n}\to\infty$. The third term goes to 00 by (3.2) and the fact that $j_{n,\tau^{\prime}}\to\infty$.
@@ 4.5. Anchored tail process
As for condition (4), $Y_{j}\geq 1$ for all $j\leq-1$ implies that the hit to $\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ at time $r_{n}$ is the first hit to that same neighbourhood of $\Lambda$ (i.e. there is no hit to $\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ before time $r_{n}$ given that there is a hit to $\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ at time $r_{n}$ recall, from (4.3), that $\{X_{r_{n}}>u_{n}(\tau)\}=G^{-r_{n}}(\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})})$). Since
@@
Let $j\geq 0$. If $G^{r_{n}}(x)\in\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ then $G^{r_{n}+j}(x)\in\Lambda_{j_{n,\tau}-j}\cup H^{(j_{n,\tau}-j)}$ provided
@@
$\begin{split}&\lim_{n\to\infty}\mathbb{P}\left\{\dfrac{1}{\tau}u_{n}^{-1}(X_{r_{n}+j})=s\>\middle|\>X_{r_{n}}>u_{n}(\tau)\right\}\\ &=\lim_{n\to\infty}\mathbb{P}\left\{\dfrac{u_{n}^{-1}(X_{r_{n}+j})}{u_{n}^{-1}(X_{r_{n}})}\dfrac{u_{n}^{-1}(X_{r_{n}})}{\tau}=s\>\middle|\>X_{r_{n}}>u_{n}(\tau)\right\}\end{split}$
@@
4.6. The $21<\alpha<$ 2 case
To prove our result in the case $21<\alpha<$ 2, we are required to satisfy the small jumps condition (3.12) and we will also need to check that the sequence $(Q_{j})_{j\in\mathbb{Z}}$, obtained from the spectral decomposition of the transformed anchored tail process, satisfies the assumption (3.13). Since we have geometric decay of $\mathcal{Q}_{j}$ (from geometric growth of $(\mathbf{\tilde{Q}}_{i,j})_{j}$), this latter condition holds in our case.
For the small jumps condition, we will now follow the proof strategy of [CNT25]. Note that they did not cover the case $\alpha=1$, so we will also omit this. As is well known (see [D83]), to show (3.12) it is sufficient to show that for $Y_{j}=X_{j}{\bf 1}_{X_{j}\leq\varepsilon a_{n}}-\mathbb{E}\left(X_{j}{\bf 1}_{X_{j}\leq\varepsilon a_{n}}\right)$, that
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{n}{a_{n}^{2}}\sum_{j=1}^{n}\max\left\{0,\mathbb{E}(Y_{1}Y_{j})\right\}=0.$
As in [CNT25], we define $\phi_{n,\varepsilon}=\phi_{\alpha}\cdot{\bf 1}_{\phi_{\alpha}\leq\varepsilon a_{n}}$ and it is sufficient to show that
[…]
@@ 4.6. The $21<\alpha<$ 2 case
[…]
@@
$\frac{n}{a_{n}^{2}}\|\phi_{n,\varepsilon}\|_{BV}\|\phi_{n,\varepsilon}\|_{L^{1}}e^{-cj}\lesssim\frac{n}{a_{n}^{2}}(a_{n}\varepsilon)^{\beta+1}e^{-\theta j}\approx n^{1+\frac{1}{\alpha}(\beta-1)}e^{-\theta j}$
@@
[…]
@@
We have shown that our observable has regularly varying tails, a well-defined anchored process, that the system satisfies the relevant dependence/mixing conditions and in the case $21<\alpha<$ 2, we have the required small jumps condition. Hence Theorem 4.1 (and hence, as in [FFT25], also Theorem 4.2) is proved.

### v2

@@ 4.4. Dependence requirements
In order to prove that the conditions $\D_{q_{n}}$ and $\D^{\prime}_{q_{n}}$ from Section 3.2 hold in our setting we will approximate the sets $A_{n,\ell}^{(q_{n})}$ by a union of cylinders of controlled depth and then use the excellent mixing properties on cylinders in the spirit of [HP14, DHY21, BF24], for example.
@@
$U_{n}(\tau^{\prime},\tau^{\prime\prime})$ is made up of holes in the dynamical construction of $\Lambda$ (i.e. subsets of the form $\Lambda_{k}\setminus\Lambda_{k+1}$ where $k\in\mathbb{N}$), where $j_{n,\tau^{\prime}}$ is as in (4.2). Moreover, as observed earlier, for $n$ sufficiently large,
$\begin{split}U_{n}(\tau^{\prime},\tau^{\prime\prime})^{(q)}&=U_{n}(\tau^{\prime},\tau^{\prime\prime})\cap G^{-1}(U_{n}(\tau^{\prime},\tau^{\prime\prime}))^{c}\cap\dots\cap G^{-q}(U_{n}(\tau^{\prime},\tau^{\prime\prime}))^{c}\\ &=U_{n}(\tau^{\prime},\tau^{\prime\prime})\cap G^{-1}(U_{n}(\tau^{\prime},\tau^{\prime\prime}))^{c}\end{split}$ (4.8)
so, in particular, taking $q_{n}=1$ for all $n\in\mathbb{N}$ is the natural choice to prove condition $\D_{q_{n}}^{{}^{\prime}*}$, which is established in 4.4.
@@
where $\beta,\gamma\in[0,1]$ and $\theta$ is as in Proposition 4.6 (as $H^{(j_{n,\tau^{\prime}})}$ is a fraction of $\Lambda_{j_{n,\tau^{\prime}}-1}\setminus\Lambda_{j_{n,\tau^{\prime}}}$ while $K$ is a fraction of $\Lambda_{j_{n,\tau^{\prime}}}\setminus\Lambda_{j_{n,\tau^{\prime}}+1}$). Observe that the larger the set $H^{(j_{n,\tau^{\prime}})}$ is, the smaller the set $K$ is and vice-versa. This means that, in particular, $\beta$ and $\gamma$ cannot be simultaneously equal to 0 or to $1$.
On the other hand, taking into consideration the definition of the function $\psi$ and of the approximating cylinders, as well as the fact that $G$ is a full branched Markov linear map (which, in particular, means that there exist $\tilde{C}>0$ and $0<\eta<1$ such that $\mathfrak{m}(\omega)\leq\tilde{C}\eta^{n}$, for all $n\in\mathbb{N}$ and $\omega\in\mathcal{P}^{n}$), we claim that
@@
As a result, we only need to check that the conditions $\D_{q_{n}}^{*}$ and $\D^{{}^{\prime}*}_{q_{n}}$ are verified for the cylinder approximating sets. To this end, we next define a strong form of mixing, which is a property of the maps we are considering here.
@@
We remark that $G$ is exponentially $\phi$-mixing, see for example [AN05, Theorem 1]. Recall that, as observed above, we only need to check the conditions $\D_{q_{n}}^{*}$, $\D_{q_{n}}^{{}^{\prime}*}$, where the set $A_{n,\ell}^{(q_{n})}$ is replaced by the approximations $\Gamma_{n}^{+}$ and $\Gamma_{n}^{-}$. In what follows, we will only check such conditions for $\Gamma_{n}^{+}$, but the argument applies with minor adjustments to $\Gamma_{n}^{-}$.
Proof of Condition $\D_{q_{n}}^{*}$.
@@
Proof of Condition $\D_{q_{n}}^{{}^{\prime}*}$.
@@
for some $\tilde{C},\hat{C}>0$. The second term goes to 0 by (3.2), Lemma 4.8 and the fact that $k_{n}\to\infty$. The third term goes to 0 by (3.2) and the fact that $j_{n,\tau^{\prime}}\to\infty$.
@@ 4.5. Anchored tail process
As for condition (4), $Y_{j}\geq 1$ for all $j\leq-1$ implies that the hit to $\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ at time $r_{n}$ is the first hit to that same neighbourhood of $\Lambda$ (i.e. there is no hit to $\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ before time $r_{n}$ given that there is a hit to $\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ at time $r_{n}$ - recall, from (4.3), that $\{X_{r_{n}}>u_{n}(\tau)\}=G^{-r_{n}}(\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})})$). Since
@@
First note that for $t\in[0,1]$, $\mathbb{P}(Y_{0}\leq t)=\mathbb{P}(F\circ\psi\leq t)=t$ by Lemma 2.2(c)(iii), so we write this uniform distribution by $U$. For $j\geq 1$, if $G^{r_{n}}(x)\in\Lambda_{j_{n,\tau}}\cup H^{(j_{n,\tau})}$ then $G^{r_{n}+j}(x)\in\Lambda_{j_{n,\tau}-j}\cup H^{(j_{n,\tau}-j)}$ provided
@@
$\begin{split}&\hskip-28.45274pt\lim_{n\to\infty}\mathbb{P}\left\{\dfrac{1}{\tau}u_{n}^{-1}(X_{r_{n}+j})=s\>\middle|\>X_{r_{n}}>u_{n}(\tau)\right\}\\ &=\lim_{n\to\infty}\mathbb{P}\left\{\dfrac{u_{n}^{-1}(X_{r_{n}+j})}{u_{n}^{-1}(X_{r_{n}})}\dfrac{u_{n}^{-1}(X_{r_{n}})}{\tau}=s\>\middle|\>X_{r_{n}}>u_{n}(\tau)\right\}\end{split}$
@@
4.6. The $1<\alpha<2$ case
To prove our result in the case $1<\alpha<2$, we are required to satisfy the small jumps condition (3.12) and we will also need to check that the sequence $(Q_{j})_{j\in\mathbb{Z}}$, obtained from the spectral decomposition of the transformed anchored tail process, satisfies the assumption (3.13). Since we have geometric decay of $\mathcal{Q}_{j}$ (from geometric growth of $(\mathbf{\tilde{Q}}_{i,j})_{j}$), this latter condition holds in our case.
For the small jumps condition, we will now follow the proof strategy of [CNT25]. Note that they did not cover the case $\alpha=1$, so we will also omit this. As is well known (see [D83]), to show (3.12) it is sufficient to show that for $Y_{j}=X_{j}{\bf 1}_{X_{j}\leq\varepsilon\mathfrak{a}_{n}}-\mathbb{E}\left(X_{j}{\bf 1}_{X_{j}\leq\varepsilon\mathfrak{a}_{n}}\right)$, that
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{n}{\mathfrak{a}_{n}^{2}}\sum_{j=1}^{n}\max\left\{0,\mathbb{E}(Y_{1}Y_{j})\right\}=0.$
As in [CNT25], we define $\phi_{n,\varepsilon}=\phi_{\alpha}\cdot{\bf 1}_{\phi_{\alpha}\leq\varepsilon\mathfrak{a}_{n}}$ and it is sufficient to show that
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{n}{\mathfrak{a}_{n}^{2}}\sum_{j=1}^{n}\int|\phi_{n,\varepsilon}|\cdot|\phi_{n,\varepsilon}\circ G^{j}|\penalty 10000\ d\mathfrak{m}=0.$
@@ 4.6. The $1<\alpha<2$ case
$\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{n}{\mathfrak{a}_{n}^{2}}\sum_{j=1}^{b\log n}\int|\phi_{n,\varepsilon}|\cdot|\phi_{n,\varepsilon}\circ G^{j}|\penalty 10000\ d\mathfrak{m}=0$
@@
$\frac{n}{\mathfrak{a}_{n}^{2}}\|\phi_{n,\varepsilon}\|_{BV}\|\phi_{n,\varepsilon}\|_{L^{1}}e^{-cj}\lesssim\frac{n}{\mathfrak{a}_{n}^{2}}(\mathfrak{a}_{n}\varepsilon)^{\beta+1}e^{-\theta j}\approx n^{1+\frac{1}{\alpha}(\beta-1)}e^{-\theta j}$
@@
For the rest of the argument, we can follow the steps of the proof of [CNT25, Theorem 8.1]. For the estimates (II) and (III) there, the argument is identical. For (I) we observe that in the setting of this paper we are always in Case 1, and the derivative is constant 3, so we get an exponentially decaying constant in front of $\int\phi_{n,\varepsilon}^{2}\penalty 10000\ d\mathfrak{m}$, so the proof then follows.
@@
We have shown that our observable has regularly varying tails, a well-defined anchored process, that the system satisfies the relevant dependence/mixing conditions and in the case $1<\alpha<2$, we have the required small jumps condition. Hence Theorem 4.1 (and hence, as in [FFT25], also Theorem 4.2) is proved.